### v1

# Yongtao Li, Lihua Feng, Yuejian Peng

Abstract
A well-known theorem of Mantel states that every $n$-vertex graph with more than $\lfloor n^{2}/4\rfloor$ edges contains a triangle. An interesting problem in extremal graph theory studies the minimum number of edges contained in triangles among graphs with a prescribed number of vertices and edges. Erdős, Faudree and Rousseau (1992) showed that a graph on $n$ vertices with more than $\lfloor n^{2}/4\rfloor$ edges contains at least $12\lfloor n/2\rfloor+$ 1 edges in triangles. Such edges are called triangular edges. In this paper, we present a spectral version of the result of Erdős, Faudree and Rousseau. Using the supersaturation-stability and the spectral technique, we prove that every $n$-vertex graph $G$ with $\lambda(G)\geq\sqrt{\lfloor n^{2}/4\rfloor}$ contains at least $12\lfloor{n}/{2}\rfloor-$ 1 triangular edges, unless $G$ is a balanced complete bipartite graph. The method in our paper has some interesting applications. Firstly, the supersaturation-stability can be used to revisit a conjecture of Erdős concerning with the booksize of a graph, which was initially proved by Edwards (unpublished), and independently by Khadžiivanov and Nikiforov (1979). Secondly, our method can improve the bound on the order $n$ of a graph by dropping the condition on $n$ being sufficiently large, which is obtained from the triangle removal lemma. Thirdly, the supersaturation-stability can be applied to deal with the spectral extremal graph problems on counting triangles, which was recently studied by Ning and Zhai (2023).
Key words: Extremal graph theory; triangular edges; spectral radius.
2010 Mathematics Subject Classification. 05C35; 05C50.
1 Introduction
Extremal combinatorics is increasingly becoming a fascinating mathematical discipline as well as an essential component of many mathematical areas, and it has experienced an impressive growth in recent years. Extremal combinatorics concerns the problems of determining the maximal or the minimal size of a combinatorial object that satisfies certain properties. One of the most important problems is the so-called Turán-type problem, which has played an important role in the development of extremal combinatorics. More precisely, the Turán-type questions usually study the maximum possible number of edges in a graph that does not contain a specific subgraph. Such kind of questions could be viewed as the cornerstone of extremal graph theory and have been studied extensively in the literature.
A graph $G$ is $F$-free if it does not contain a subgraph isomorphic to $F$. For example, every bipartite graph is triangle-free. A classical result in extremal graph theory is Mantel’s theorem [5], which asserts that every triangle-free graph on $n$ vertices contains at most $\lfloor n^{2}/4\rfloor$ edges. This result is tight by considering the bipartite Turán graph $T_{n,2}$, where $T_{n,2}$ is a complete bipartite graph whose two vertex parts have sizes as equal as possible. Equivalently, each graph on $n$ vertices with more than $\lfloor n^{2}/4\rfloor$ edges must contain a triangle.
There are several results in the literature that guarantee something stronger than just one triangle. For example, in 1941, Rademacher (unpublished paper, see Erdős [15, 18]) showed that such graphs contain at least $\lfloor n/2\rfloor$ triangles. After this result, Erdős [16, 17] showed that there exists a small constant $c>0$ such that if $n$ is large enough and $1\leq q<cn$, then every $n$-vertex graph with $\lfloor n^{2}/4\rfloor+q$ edges has at least $q\lfloor{n}/{2}\rfloor$ triangles. Furthermore, Erdős conjectured the constant $c={1}/{2}$, which was finally confirmed by Lovász and Simonovits [39, 40] in 1975, they proved that if $1\leq q<{n}/{2}$ is a positive integer and $G$ is an $n$-vertex graph with $e(G)\geq\lfloor{n^{2}}/{4}\rfloor+q$, then $G$ contains at least $q\lfloor{n}/{2}\rfloor$ triangles. We refer the readers to [62, 38, 4] for recent generalizations on the Erdős–Rademacher problem. Moreover, Lovász and Simonovits [40] also studied the supersaturation problem for cliques in the case $q=o(n^{2})$. While for $q=\Omega(n^{2})$, this problem turns out to be notoriously difficult. Some recent progresses were presented by Razborov [54], Nikiforov [49], Reiher [55], Liu, Pikhurko and Staden [37]. In addition, the supersaturation problems for color-critical graphs were studied by Mubayi [42], and Pikhurko and Yilma [53].
1.1 Minimizing the number of triangular edges
In this paper, we shall consider the supersaturation problem from a different point of view. An edge is called triangular if it is contained in a triangle. We shall consider the problem on counting the number of triangular edges, rather than the number of triangles. The first result in this topic was obtained by Erdős, Faudree and Rousseau [19], who provided a tight bound on the number of triangular edges in any $n$-vertex graph with more than $\lfloor n^{2}/4\rfloor$ edges.
Theorem 1.1 (Erdős–Faudree–Rousseau, 1992).
Let $G$ be a graph with $n$ vertices and
$e(G)>e(T_{n,2}).$
Then $G$ has at least $12\lfloor\frac{n}{2}\rfloor+$ 1 triangular edges.
This bound is the best possible simply by adding an edge to the larger vertex part of the balanced complete bipartite graph. Motivated by the problem about the number of triangles, it is natural to ask how many triangular edges an $n$-vertex graph with $m$ edges must have, where $m$ is an integer satisfying $\lfloor n^{2}/4\rfloor<m\leq{n\choose 2}$. Indeed, this problem was recently studied by Füredi and Maleki [23] as well as Gruslys and Letzter [24]. Given integers $a,b,c$, let $G(a,b,c)$ denote the graph on $n=a+b+c$ vertices, which consists of a clique $A$ of size $a$ and two independent sets $B$ and $C$ of sizes $b$ and $c$ respectively, such that all edges between $B$ and $A\cup C$ induces a complete bipartite graph $K_{b,a+c}$. In other words, the graph $G(a,b,c)$ can be obtained from $K_{b,a+c}$ by embedding a clique of order $a$ into the part of size $a+c$. Note that $G(a,b,c)$ has (${a\choose 2}+(a+c)b$ edges and it has (${a\choose 2}+ab=m-bc$ triangular edges. In 2017, Füredi and Maleki [23] conjectured that the minimizers of the number of triangular edges are graphs of the form $G(a,b,c)$ or subgraphs of such graphs.
Conjecture 1.2 (Füredi–Maleki, 2017).
Let $m>n^{2}/4$ and $G$ be an $n$-vertex graph with $m$ edges that minimizes the number of triangular edges. Then $G$ is isomorphic to a subgraph of $G(a,b,c)$ for some $a,b,c$.
Particularly, Füredi and Maleki [23] proposed a numerical conjecture, which states that every $n$-vertex graph with $m$ edges has at least $g(n,m)$ triangular edges, where
$g(n,m)=\min\left\{m-bc:a+b+c=n,{a\choose 2}+b(a+c)\geq m\right\}.$
We remark that Conjecture 1.2 characterizes the structures of the minimizers, while the latter conjecture gives a lower bound only. By using a generalization of Zykov’s symmetrization method, Füredi and Maleki [23] showed an asymptotic formula: if $G$ is a graph on $n$ vertices with $m>n^{2}/4$ edges, then $G$ has at least $g(n,m)-3n/2$ triangular edges. Soon after, Gruslys and Letzter [24] proved an exact version of the result of Füredi and Maleki. Let $\mathbf{NT}(G)$ be the set of non-triangular edges of $G$. The following result was established in [24].
Theorem 1.3 (Gruslys–Letzter, 2018).
There is $n_{0}$ such that for any graph $G$ on $n\geq n_{0}$ vertices, there exists a graph $H=G(a,b,c)$ on $n$ vertices such that $e(H)\geq e(G)$ and $|\mathbf{NT}(H)|\geq|\mathbf{NT}(G)|$.
Theorem 1.3 shows that for sufficient large $n$, the minimum number of triangular edges among all $n$-vertex graphs with at least $m$ edges is attained by the graph $G(a,b,c)$ or its subgraph for some $a,b,c$. We refer the readers to [24] for more details and [25] for the study on the minimum number of edges that occur in odd cycles.
1.2 Spectral extremal graph problems
Spectral graph theory aims to apply the eigenvalues of matrices associated with graphs to find the structural information of graphs. Let $G$ be a simple graph on vertex set $\{v_{1},v_{2},\ldots,v_{n}\}$. The adjacency matrix of $G$ is defined as $A(G)=[a_{i,j}]_{i,j=1}^{n}$, where $a_{i,j}=1$ if $v_{i}$ and $v_{j}$ are adjacent, and $a_{i,j}=0$ otherwise. Let $\lambda(G)$ be the spectral radius of $G$, which is defined as the maximum of modulus of eigenvalues of $A(G)$. The study in this article mainly concentrates on the adjacency spectral radius.
As mentioned before, the Turán type problem studies the maximal size of a graph forbidding certain subgraphs. In particular, one could wish to investigate the maximum possible spectral radius of the associated adjacency matrix of a graph which does not contain certain subgraphs. The interplay between these two areas above is called the spectral Turán-type problem. One of the famous results of this type was obtained in 1986 by Wilf [61] who showed that every graph $G$ on $n$ vertices with $\lambda(G)>(1-{1}/{r})n$ contains a clique $K_{r+1}$. This spectral version generalized the classical Turán theorem by invoking the fact $\lambda(G)\geq 2m/n$. It is worth emphasizing that spectral Turán problems have been receiving considerable attention in the last two decades and it is still an attractive topic; see, e.g., [61, 45, 46, 30] for graphs with no cliques, [7, 35, 67, 14] for a conjecture of Bollobás and Nikiforov, [35, 64, 33] for non-bipartite triangle-free graphs, [58, 34] for planar graphs and outplanar graphs, [47] for a spectral Erdős–Stone–Bollobás theorem, [48] for the spectral stability theorem, [10, 28] for spectral problems on cycles, [9] for a spectral Erdős–Sós theorem, [20] for some specific trees, [63] for a spectral Erdős–Pósa theorem, [66, 50] for books and theta graphs, [27, 68] for cycles of consecutive lengths, [60] for a spectral conjecture on a class of graphs, and [59, 65] for graphs without $K_{t}$-minors or $K_{s,t}$-minors.
Although there has been a wealth of research on the spectral extremal problems in recent years, there are very few conclusions on problems of counting substructures in terms of the spectral radius. The first result in this topic can even be traced back to a result of Bollobás and Nikiforov [7] who showed that for every $n$-vertex graph $G$ and $r\geq 2$, the number of cliques of order $r+1$ satisfies
$k_{r+1}(G)\geq\left(\frac{\lambda(G)}{n}-1+\frac{1}{r}\right)\frac{r(r-1)}{r+1}\left(\frac{n}{r}\right)^{r+1}.$
This result provides a spectral version of supersaturation on cliques.
In 2023, Ning and Zhai [52] studied the spectral saturation on triangles. A result of Erdős and Rademacher states that every $n$-vertex graph $G$ with $e(G)>e(T_{n,2})$ contains at least $\lfloor\frac{n}{2}\rfloor$ triangles. Correspondingly, it is natural to consider the spectral version: if $G$ is a graph with $\lambda(G)>\lambda(T_{n,2})$, whether $G$ has at least $\lfloor\frac{n}{2}\rfloor$ triangles. Unfortunately, this result is not true. Let $K_{a,b}^{+}$ be the graph obtained from the complete bipartite graph $K_{a,b}$ by adding an edge to the vertex set of size $a$. For even $n$, we take $a=\frac{n}{2}+1$ and $b=\frac{n}{2}-1$. One can verify that $\lambda(K_{\frac{n}{2}+1,\frac{n}{2}-1}^{+})>\lambda(T_{n,2})$, while $K_{\frac{n}{2}+1,\frac{n}{2}-1}^{+}$ has exactly $\frac{n}{2}-1$ triangles. Recently, Ning and Zhai [52] provided the following tight bound.
Theorem 1.4 (Ning–Zhai, 2023).
If $G$ is an $n$-vertex graph with
$\lambda(G)\geq\lambda(T_{n,2}),$
then $G$ has at least $\left\lfloor\frac{n}{2}\right\rfloor-1$ triangles, unless $G$ is the bipartite Turán graph $T_{n,2}$.
2 Main results
In the sequel, we shall put our attention on the extremal graph problems concerning the spectral supersaturation. Specifically, we shall present a tight bound on the number of triangular edges in a graph with spectral radius larger than that of $T_{n,2}$. Hence, we prove a spectral version of the result of Erdős, Faudree and Rousseau.
Theorem 2.1.
Let $G$ be a graph with $n\geq 5432$ vertices and
$\lambda(G)\geq\lambda(T_{n,2}).$
Then $G$ has at least $12\lfloor\frac{n}{2}\rfloor-$ 1 triangular edges, unless $G=T_{n,2}$.
The spectral condition in Theorem 2.1 is easier to satisfy than the edge-condition in Theorem 1.1. Namely, if a graph $G$ satisfies $e(G)>e(T_{n,2})$, then $\lambda(G)>\lambda(T_{n,2})$. This observation can be guaranteed by $\lambda(G)\geq 2e(G)/n$. Nevertheless, there are many graphs with $\lambda(G)>\lambda(T_{n,2})$ but $e(G)<e(T_{n,2})$. Let $S_{n,k}$ be the split graph, which is the join of a clique of size $k$ with an independent set of size $n-k$. Taking $k=n/5$, we can verify that $S_{n,k}$ is a required example. A few words regarding the tightness of Theorem 2.1 are due. We show in next section that there exist three graphs $G$ such that $\lambda(G)>\lambda(T_{n,2})$ and $G$ has exactly $12\lfloor\frac{n}{2}\rfloor-$ 1 triangular edges, which implies the bound in Theorem 2.1 is tight.
It is reasonable to reach such a difference between the results on the spectral radius and the size of a graph. Note that if $e(G)>e(T_{n,2})$, then $e(G)\geq e(T_{n,2})+1$ holds immediately. While, if $\lambda(G)>\lambda(T_{n,2})$ holds, then there are many graphs with $\lambda(G)$ very closing to $\lambda(T_{n,2})$ and $e(G)=e(T_{n,2})$. Roughly speaking, the spectral radii of graphs are distributed more compactly. Motivated by this observation, Li, Lu and Peng [32] proposed a spectral conjecture on Mubayi’s result [42] and showed a spectral version of Erdős–Rademacher’s theorem. Next, we are going to provide a variant of Theorem 2.1. We shall establish a spectral condition corresponding to the edge condition $e(G)\geq e(T_{n,2})+1$. Recall that $K_{\lceil\frac{n}{2}\rceil,\lfloor\frac{n}{2}\rfloor}^{+}$ is the graph obtained from the complete bipartite graph $K_{\lceil\frac{n}{2}\rceil,\lfloor\frac{n}{2}\rfloor}$ by adding an edge to the vertex part of size $\lceil\frac{n}{2}\rceil$.
Theorem 2.2.
Let $G$ be a graph on $n\geq 5432$ vertices with
$\lambda(G)\geq\lambda(K_{\lceil\frac{n}{2}\rceil,\lfloor\frac{n}{2}\rfloor}^{+}).$
Then $G$ has at least $12\lfloor\frac{n}{2}\rfloor+$ 1 triangular edges, with equality if and only if $G=K_{\lceil\frac{n}{2}\rceil,\lfloor\frac{n}{2}\rfloor}^{+}$.
Our proofs of Theorems 2.1 and 2.2 are quite different from that of Theorem 1.4. It is a classical spectral method to use the Perron eigenvector together with the walks of length two to deduce the structural properties of spectral extremal graphs; see, e.g., [58, 59, 9, 34, 52, 65]. However, applying this spectral method turns out to be difficult for graphs with much more triangles or triangular edges. The key ingredient in our proof attributes to a supersaturation-stability result (Theorem 4.5), which says roughly that if a graph is far from being bipartite, then it contains a large number of triangles. This result may be of independent interest. Although we used the stability method, we only need a weak bound $n\geq 5432$ exactly***It seems possible to obtain a slightly better bound. To avoid unnecessary and tedious calculations, we did not attempt to get the best bound on the order of the graph in our proof. , instead of the strong condition that $n$ is sufficiently large. Apart from the supersaturation-stability, another technique used in this paper is a spectral technique developed by Cioabă, Feng, Tait and Zhang [8]; see, e.g., [29, 13, 60] for recent results. Furthermore, we will obtain some approximately structural results that describe the almost-extremal graphs with large spectral radius and few triangular edges.
We anticipate that with additional efforts, this method used in the proof of Theorems 2.1 and 2.2 could be possibly applied to treat some spectral extremal graph problems in which the desired extremal graph contains a small number of triangles. Incidentally, an upper bound on the number of triangular edges eventually leads to a restriction on the number of triangles. In particular, we shall present three quick applications of our method in Subsection 4.3. The first application gives a short proof of a conjecture of Erdős, which asserts that every $n$-vertex graph with more than $n^{2}/4$ edges contains more than $n/6$ triangles sharing a common edge. The second application allows us to simplify the proof of the main result in [8], and it can also improve the bound on the order of graphs, which was previously obtained from the celebrated triangle removal lemma. The last application provides an alternative proof for the elegant result of Ning and Zhai stated in Theorem 1.4.
Organization. In Section 3, we shall present some computations on the spectral radius of the expected extremal graphs. As mentioned before, these graphs reveal that the bound in Theorem 2.1 can not be improved. Moreover, we will show the spectral version of the supersaturation for triangular edges (Theorems 3.5 and 3.7). One of the key ideas in this paper, i.e., the supersaturation-stability (Theorems 4.3 and 4.5), would be introduced in Section 4. As indicated above, some quick applications of the supersaturation-stability method will be provided in Subsection 4.3. In Sections 5 and 6, we will present the detailed proofs of Theorems 2.1 and 2.2, respectively. After proving our results, we propose some related spectral extremal problems involving the edges that occur in cliques or odd cycles.
Notation. We usually write $G=(V,E)$ for a graph with vertex set $V=\{v_{1},\ldots,v_{n}\}$ and edge set $E=\{e_{1},\ldots,e_{m}\}$, where we admit $n=|V|$ and $m=|E|$. If $S\subseteq V$ is a subset of the vertex set, then $G[S]$ denotes the subgraph of $G$ induced by $S$, i.e., the graph on $S$ whose edges are those edges of $G$ with both endpoints in $S$. By convention, we denote $e(S)=e(G[S])$. We will write $G[S,T]$ for the induced subgraph of $G$ whose edges have one endpoint in $S$ and the other in $T$, and similarly, we write $e(S,T)$ for the number of edges of $G[S,T]$. Let $N(v)$ be the set of vertices adjacent to a vertex $v$ and let $d(v)=|N(v)|$. Moreover, we denote $N_{S}(v)=N(v)\cap S$ and $d_{S}(v)=|N_{S}(v)|$ for simplicity. We will write $t(G)$ for the number of triangles of $G$. For an integer $p\geq 3$, we write $k_{p}(G)$ for the number of cliques of order $p$ in $G$.
3 Some auxiliary lemmas
3.1 Computations for extremal graphs
We will show that Theorem 2.1 is the best possible. Recall that $K_{a,b}^{+}$ denotes the graph obtained from a complete bipartite graph $K_{a,b}$ by adding an edge to the vertex part of size $a$. The following three graphs have spectral radii larger than $\lambda(T_{n,2})$ and they contain exactly $12\lfloor n/2\rfloor-$ 1 triangular edges. Moreover, these graphs have exactly $\lfloor n^{2}/4\rfloor$ edges.
Figure 1: The graphs $K_{\frac{n}{2}+1,\frac{n}{2}-1}^{+},K_{\frac{n}{2},\frac{n}{2}}^{+|}$ and $K_{\frac{n+1}{2},\frac{n-1}{2}}^{+|}$.
Lemma 3.1.
If $n\geq 4$ is even, then
$\lambda(K_{\frac{n}{2}+1,\frac{n}{2}-1}^{+})>\lambda(T_{n,2}).$
Proof.
Upon computation, it follows that $\lambda(K_{\frac{n}{2}+1,\frac{n}{2}-1}^{+})$ is the largest root of
$f_{1}(x)=x^{3}-x^{2}+x-{(n^{2}x)}/{4}+{n^{2}}/{4}-n+1.$
Since $f_{1}(\frac{n}{2})=1-\frac{n}{2}<0$ for $n\geq 4$, we have $\lambda(K_{\frac{n}{2}+1,\frac{n}{2}-1}^{+})>\lambda(T_{n,2})=\frac{n}{2}$. ∎
Lemma 3.2.
Let $G=K_{\frac{n}{2},\frac{n}{2}}^{+|}$ be the graph obtained from $K_{\frac{n}{2},\frac{n}{2}}$ by adding an edge $e_{1}$ into the part of size $\frac{n}{2}$ and deleting an edge $e_{2}$ between two parts such that $e_{2}$ is incident to $e_{1}$. Then
$\lambda(K_{\frac{n}{2},\frac{n}{2}}^{+|})>\lambda(T_{n,2}).$
Proof.
By calculation, we obtain that $\lambda(K_{\frac{n}{2},\frac{n}{2}}^{+|})$ is the largest root of
$f_{2}(x)=x^{4}-{(n^{2}x^{2})}/{4}-(n-2)x+1+{n^{2}}/{2}-2n.$
One can check that $f_{2}(\frac{n}{2})=1-n<0$ and hence $\lambda(K_{\frac{n}{2},\frac{n}{2}}^{+|})>\frac{n}{2}=\lambda(T_{n,2})$. ∎
Lemma 3.3.
If $n\geq 5$ is odd and $G=K_{\frac{n+1}{2},\frac{n-1}{2}}^{+|}$ is the graph obtained from $K_{\frac{n+1}{2},\frac{n-1}{2}}$ by adding an edge $e_{1}$ into the part of size $\frac{n+1}{2}$ and deleting an edge $e_{2}$ between two parts such that $e_{2}$ is incident to $e_{1}$, then
$\lambda(K_{\frac{n+1}{2},\frac{n-1}{2}}^{+|})>\lambda(T_{n,2}).$
Proof.
By direct calculation, we know that $\lambda(K_{\frac{n+1}{2},\frac{n-1}{2}}^{+|})$ is the largest root of
$f_{3}(x)=x^{4}-{(n^{2}x^{2})}/{4}+x^{2}/4-(n-3)x+{n^{2}}/{2}-2n+{3}/{2}.$
We can verify that
$f_{3}\Bigl{(}\frac{1}{2}\sqrt{n^{2}-1}\Bigr{)}=\frac{1}{2}(n-3)\left(n-1-\sqrt{n^{2}-1}\right)<0,$
which implies $\lambda(K_{\frac{n+1}{2},\frac{n-1}{2}}^{+|})>\frac{1}{2}\sqrt{n^{2}-1}=\lambda(T_{n,2})$, as desired. ∎
The following lemma will be used in the proof of Theorem 2.2, and it provides a characterization of the spectral radius of the graph $K_{\lceil\frac{n}{2}\rceil,\lfloor\frac{n}{2}\rfloor}^{+}$.
Lemma 3.4.
(a) If $n$ is even, then $\lambda(K_{\frac{n}{2},\frac{n}{2}}^{+})$ is the largest root of
$f(x)=x^{3}-x^{2}-(n^{2}x)/4+n^{2}/4-n.$
(b) If $n$ is odd, then $\lambda(K_{\frac{n+1}{2},\frac{n-1}{2}}^{+})$ is the largest root of
$g(x)=x^{3}-x^{2}+x/4-(n^{2}x)/4+n^{2}/4-n+3/4.$
Consequently, for $n\geq 4$, we have
$\lambda^{2}(K_{\lceil\frac{n}{2}\rceil,\lfloor\frac{n}{2}\rfloor}^{+})>\left\lfloor{n^{2}}/{4}\right\rfloor+2.$
Proof.
By calculation, we can verify that for even $n$,
$f(\sqrt{{n^{2}}/{4}+2})=\sqrt{n^{2}+8}-n-2<0,$
and for every odd $n$,
$g(\sqrt{{(n^{2}-1)}/{4}+2})=\sqrt{n^{2}+7}-n-1<0.$
So we get $\sqrt{\left\lfloor{n^{2}}/{4}\right\rfloor+2}<\lambda(K_{\lceil\frac{n}{2}\rceil,\lfloor\frac{n}{2}\rfloor}^{+})$. This completes the proof. ∎
3.2 Spectral supersaturation for triangular edges
Recall that $t(G)$ denotes the number of triangles in a graph $G$. A special case of an aforementioned result of Bollobás and Nikiforov [7] states that
$t(G)\geq\frac{n^{2}}{12}\left(\lambda-\frac{n}{2}\right).$
From this inequality, we can obtain a spectral supersaturation for triangular edges. We denote by $\lambda(G)/n$ the spectral density of a graph $G$. Informally, once the spectral density of a graph exceeds that of the bipartite Turán graph, we not only find $12\lfloor\frac{n}{2}\rfloor-$ 1 triangular edges, but in fact a large number of triangular edges with positive density, i.e., there are $\Omega(n^{2})$ triangular edges. This gives a phase transition type result.
Theorem 3.5.
If $\varepsilon>0$ and $G$ is a graph on $n$ vertices with
$\lambda(G)\geq\frac{n}{2}+\varepsilon n,$
then $G$ contains $32^{-1/3}\varepsilon^{2/3}n^{2}$ triangular edges.
Proof.
First of all, it follows from $\lambda\geq\frac{n}{2}+\varepsilon n$ that
$t(G)\geq\frac{n^{2}}{12}\left(\lambda-\frac{n}{2}\right)\geq\frac{\varepsilon}{12}n^{3}.$
Let $m^{\prime}$ be the number of triangular edges of $G$, and let $G^{\prime}$ be the subgraph of $G$ whose edges consist of all triangular edges of $G$. Clearly, we have $t(G)=t(G^{\prime})$. Suppose on the contrary that $m^{\prime}\leq 32^{-1/3}\varepsilon^{2/3}n^{2}$. Then applying Kruskal–Katona’s theorem (see, e.g., [5, page 305]) on $G^{\prime}$ gives $t(G^{\prime})\leq{\sqrt{2m^{\prime}}\choose 3}<\frac{\varepsilon}{12}n^{3}$, a contradiction. Therefore, we have $m^{\prime}>32^{-1/3}\varepsilon^{2/3}n^{2}$, so $G$ has more than $32^{-1/3}\varepsilon^{2/3}n^{2}$ triangular edges. ∎
In our proofs of Theorems 2.1 and 2.2, we need to use the following lemma, which counts the number of triangles in terms of the spectral radius and the size of a graph.
Lemma 3.6 (See [7, 8, 52]).
Let $G$ be a graph with $m$ edges. Then
$t(G)\geq\frac{\lambda\bigl{(}\lambda^{2}-m\bigr{)}}{3}.$
The equality holds if and only if $G$ is a complete bipartite graph.
The inequality can be written as the following versions:
$\lambda^{3}\leq 3t+m\lambda\quad\Leftrightarrow\quad{t\geq\frac{1}{3}\lambda(\lambda^{2}-m)}\quad\Leftrightarrow\quad m\geq\lambda^{2}-\frac{3t}{\lambda}.$
This inequality was firstly published by Bollobás and Nikiforov as a special case of their result [7, Theorem 1], and it was independently proved by Cioabă, Feng, Tait and Zhang [8], and also proved by Ning and Zhai [52]. These three proofs are quite different. Moreover, as demonstrated by Ning and Zhai [52], the equality in Lemma 3.6 holds if and only if $G$ is a complete bipartite graph (possibly with some isolated vertices). Indeed, this result could be seen from $t(G)=\frac{1}{6}\sum_{i=1}^{n}\lambda_{i}^{3}$, where $\lambda_{1}\geq\lambda_{2}\geq\cdots\geq\lambda_{n}$ are eigenvalues of $G$. Then $t(G)-\frac{1}{3}\lambda_{1}(\lambda_{1}^{2}-m)=\frac{1}{6}\sum_{i=2}^{n}(\lambda_{1}+\lambda_{i})\lambda_{i}^{2}\geq 0$.
From Lemma 3.6, we can see that every graph with $\lambda(G)>\sqrt{m}$ contains a triangle. Next, we show a spectral supersaturation on the number of triangular edges.
Theorem 3.7.
If $\varepsilon>0$ and $G$ is a graph with $m$ edges and
$\lambda(G)\geq(1+\varepsilon)\sqrt{m},$
then $G$ contains $2^{1/3}\varepsilon^{2/3}m$ triangular edges.
Proof.
Since $\lambda\geq(1+\varepsilon)\sqrt{m}$, Lemma 3.6 implies
$t(G)\geq\frac{\lambda(\lambda^{2}-m)}{3}>\frac{2\varepsilon}{3}m^{3/2}.$
Let $G^{\prime}$ be the subgraph in $G$ whose edges consist of all triangular edges of $G$. Suppose on the contrary that $G^{\prime}$ has $m^{\prime}\leq 2^{1/3}\varepsilon^{2/3}m$ edges. Then by Kruskal–Katona’s theorem (see, e.g., [5, page 305]), we get $t(G^{\prime})\leq{\sqrt{2m^{\prime}}\choose 3}<\frac{2}{3}\varepsilon m^{3/2}$, a contradiction. Therefore, we have $m^{\prime}>2^{1/3}\varepsilon^{2/3}m$, and $G$ has more than $2^{1/3}\varepsilon^{2/3}m$ triangular edges. ∎
4 The supersaturation-stability results
4.1 The Lovász–Simonovits stability
To prove and generalize Erdős’ conjecture on triangle-supersaturated graphs, Lovász and Simonovits [39] proved a stability result, and a much more general theorem in [40], the simplest form of which is the following:
Theorem 4.1 (Lovász–Simonovits stability).
For any constant $C>0$, there exists an $\varepsilon>0$ such that if $|k|<\varepsilon n^{2}$ and $G$ is an $n$-vertex graph with $\lfloor n^{2}/4\rfloor+k$ edges and fewer than $C|k|n$ triangles, then one can remove $O(|k|)$ edges from $G$ to get a bipartite graph.
It was shown in [40] that if $G$ is an $n$-vertex graph with $e(G)=(1-\frac{1}{x})\frac{n^{2}}{2}$ edges, where $x>1$ is a real number, then for any integer $p\leq x+1$, the number of $p$-cliques satisfies $k_{p}(G)\geq{x\choose p}(\frac{n}{x})^{p}$; see, e.g., [41, p. 449] for a detailed proof. In the next, we introduce a more general theorem on stability. Let $T_{n,p}$ denote the $p$-partite Turán graph on $n$ vertices, that is, $T_{n,p}$ is a complete $p$-partite graph whose parts have sizes as equal as possible.
Theorem 4.2 (Lovász–Simonovits, 1983).
Let $C>0$ be an arbitrary constant. There exist constants $\delta>0$ and $C^{\prime}>0$ such that if $1\leq k<\delta n^{2}$ and $G$ is an $n$-vertex graph with $e(G)=(1-\frac{1}{x})\frac{n^{2}}{2}$, and $p\leq x+1$ is an integer satisfying $e(G)=e(T_{n,p})+k$ and
$k_{p}(G)<{x\choose p}\left(\frac{n}{x}\right)^{p}+Ckn^{p-2},$
then $G$ can be made $\lfloor x\rfloor$-partite by removing at most $C^{\prime}k$ edges.
The application of the Lovász–Simonovits stability can be replaced here by an easy application of the graph removal lemma [11] and the Erdős–Simonovits stability [56]. The former result was initially proved using the Szemerédi Regularity Lemma, i.e., the graph regularity method. For completeness, we present the following supersaturation-stability theorem. In addition, we refer the readers to [1, 12, 21] for some similar applications on extremal set theory and Ramsey theory.
Theorem 4.3.
For any $\varepsilon>0$ and $r\geq 2$, there exist $\eta>0,\delta>0$ and $n_{0}\in\mathbb{N}$ such that if $G$ is a graph on $n\geq n_{0}$ vertices with at most $\eta n^{r+1}$ copies of $K_{r+1}$ and
$e(G)\geq\left(1-\frac{1}{r}-\delta\right)\frac{n^{2}}{2},$
then $G$ can be made $r$-partite by removing at most $\varepsilon n^{2}$ edges.
Proof.
The graph removal lemma allows us to pass to a $K_{r+1}$-free subgraph $G^{\prime}$ of $G$ which still has very many edges. At this point, we can apply the standard stability theorem to deduce that $G^{\prime}$ is nearly $r$-partite. Since we deleted few edges to go from $G$ to $G^{\prime}$, we must also have that $G$ is nearly $r$-partite. ∎
Although such an analogue can easily be obtained via the graph removal lemma, this gives bounds which are far from sufficient for our purposes. However, this approach would be less elementary and effective. In next subsection, we shall give a more efficient stability result so that we can calculate some explicit constants.
4.2 A generalized Moon–Moser’s inequality
First of all, we shall present a result of Moon and Moser [43], which counts the minimum number of triangles in a graph with given order and size. The alternative proofs can also be found in [5, p. 297] and [41, p. 443].
Theorem 4.4 (Moon–Moser, 1962).
Let $G$ be a graph on $n$ vertices with $m$ edges. Then
$t(G)\geq\frac{4m}{3n}\left(m-\frac{n^{2}}{4}\right),$
where the equality holds if and only if $m=e(T_{n,r})$ with $r$ dividing $n$.
We illustrate that Moon–Moser’s theorem implies a supersaturation on triangles for graph with more than $n^{2}/4$ edges. For example, if $G$ has at least ${n^{2}}/{4}+1$ edges, then it contains at least ${n}/{3}$ triangles. This result is slightly weaker than Erdős–Rademacher’s theorem. Moreover, Moon–Moser’s theorem yields that if $\varepsilon>0$ and $G$ has at least ${n^{2}}/{4}+\varepsilon n^{2}$ edges, then $G$ contains more than ${\varepsilon}n^{3}/3$ triangles. In what follows, we shall show a generalization for graphs with less than $n^{2}/4$ edges.
We say that a graph $G$ is $t$-far from being bipartite if $G^{\prime}$ is not bipartite for every subgraph $G^{\prime}$ of $G$ with $e(G^{\prime})>e(G)-t$. In other words, if $G$ is $t$-far from being bipartite, then no matter how we delete less than $t$ edges from $G$, the resulting graph is not bipartite. Equivalently, we must remove at least $t$ edges from $G$ to make it being bipartite. It is well-known that every graph $G$ contains a bipartite subgraph $H$ with $e(H)\geq e(G)/2$. From this observation, we know that if $G$ is said to be $t$-far from being bipartite, then we always admit the natural condition $t\leq e(G)/2$.
Next, we present a counting result, which comes from the work of Balogh, Bushaw, Collares, Liu, Morris and Sharifzadeh [3] during the study on the typical structure of graphs with no large cliques. This result allows us to avoid the use of the triangle removal lemma, so that we could obtain a better bound on the order of the extremal graphs. This will be explained in the forthcoming Subsection 4.3.
Theorem 4.5 (See [3]).
If $G$ is $t$-far from being bipartite, then
$t(G)\geq\frac{n}{6}\left(m+t-\frac{n^{2}}{4}\right).$
For completeness, we provide a detailed proof for readers. This result can be proved by applying a similar argument due to Sudakov [57] and Füredi [22] as well.
Proof.
We may assume that $t\leq m/2$ and $m\geq{n^{2}}/{4}-t$. This leads to $m\geq{n^{2}}/{6}$. For each $v\in V(G)$, we denote $N_{v}=N(v)$ and $N_{v}^{c}=V(G)\setminus N(v)$. Since $G$ is $t$-far from being bipartite, it follows that for every $v\in V(G)$,
$e(N_{v})+e(N_{v}^{c})\geq t.$
On the one hand, we have
$\sum_{w\in N_{v}^{c}}d(w)=2e(N_{v}^{c})+e(N_{v}^{c},N_{v})=e(N_{v}^{c})+m-e(N_{v})\geq m+t-2e(N_{v}).$
Summing over all vertices $v\in V(G)$ yields
$\sum_{v\in V(G)}\sum_{w\in N_{v}^{c}}d(w)\geq mn+nt-2\sum_{v\in V(G)}e(N_{v})=mn+nt-6t(G),$
where we used the fact $\sum_{v\in V(G)}e(N_{v})=3t(G)$. On the other hand, we get
$\sum_{v\in V(G)}\sum_{w\in N_{v}^{c}}d(w)=\sum_{v\in V(G)}\left(2m-\sum_{w\in N_{v}}d(w)\right)=2mn-\sum_{w\in V(G)}d^{2}(w).$
Combining these two inequalities, we obtain
$6t(G)\geq nt-mn+\sum_{w\in V(G)}d^{2}(w)\geq nt-mn+\frac{4m^{2}}{n}.$
Observe that ${4m^{2}}/{n}\geq 2mn-{n^{3}}/{4}$. The required bound holds immediately. ∎
The following result is a direct consequence of Theorem 4.5.
Corollary 4.6.
If $G$ is a graph with $m={n^{2}}/{4}-q$ edges, where $q\in\mathbb{Z}$, and $G$ has at most $t$ triangles, then we can remove at most ${6t}/{n}+q$ edges to make it bipartite, so $G$ has a bipartite subgraph with size at least ${n^{2}}/{4}-{6t}/{n}-2q$.
This corollary can also be deduced from a result of Sudakov [57, Lemma 2.3].
4.3 Applications of the supersaturation-stability
As promised, we shall present three quick applications of the supersaturation-stability method, i.e., the main ingredient stated in Theorem 4.5.
① Recall that a book of size $t$ consists of $t$ triangles sharing a common edge. The study of bounding the largest size of a book in a graph was initially investigated by Erdős [16] who proved that every $n$-vertex graph with at least $\lfloor n^{2}/4\rfloor+1$ edges contains a book of size $n/6-O(1)$, and he conjectured that the $O(1)$ term can be removed. This conjecture was later proved by Edwards (unpublished, see [19, Lemma 4]) and independently by Khadžiivanov and Nikiforov [26]. Here, we show that Theorem 4.5 can easily confirm the Erdős conjecture. More precisely, we can use Theorem 4.5 to prove that every graph $G$ on $n$ vertices with more than $n^{2}/4$ edges contains a book of size greater than $n/6$. Indeed, assume that $G$ has exactly $t$ triangles, then Theorem 4.5 yields that $G$ is not ${6t}/{n}$-far from being bipartite. Specifically, one can remove less than ${6t}/{n}$ edges from $G$ to destroy all $t$ triangles. So one of these edges must be contained in more than $n/6$ triangles, as needed. For more related results, we refer the readers to [6, 50, 66] and the references therein.
② In 2020, Cioabă, Feng, Tait and Zhang [8] studied the spectral extremal graphs of order $n$ for the friendship graph $F_{k}$ and sufficiently large $n$, where $F_{k}$ is the graph that consists of $k$ triangles sharing a vertex. While their proof uses the Ruzsa–Szemerédi triangle removal lemma, which leads to a worse bound on the order of the extremal graphs. Actually, the removal lemma settles the problem in the case where $k$ is fixed, and the result is meaningless when $k$ is large and growth with $n$ (say, when $k\geq\log n$). By applying Theorem 4.5, instead of the triangle removal lemma, we can not only simplify the original proof by proving the following lemma, but also show that the main result in [8] still holds for every $n\geq Ck^{4}$ with an absolute constant $C>0$.
Lemma 4.7.
If $G$ is an $F_{k}$-free graph on $n$ vertices and $\lambda(G)\geq{n}/{2}$, then
$e(G)>\frac{n^{2}}{4}-54k^{2}$
and there exists a vertex partition of $G$ as $V(G)=S\cup T$ such that
$e(S)+e(T)<108k^{2}.$
Moreover, we have
$\frac{n}{2}-13k<|S|,|T|<\frac{n}{2}+13k$
and
$\frac{n}{2}-56k^{2}<\delta(G)\leq\lambda(G)\leq\Delta(G)<\frac{n}{2}+14k.$
Proof.
A result due to Alon and Shikhelman [2, Lemma 3.1] states that if $G$ is $F_{k}$-free, then $G$ has less than $(9k-15)(k+1)n<9k^{2}n$ triangles. Using Lemma 3.6, we have $e(G)\geq\lambda^{2}-{(3t)}/{\lambda}\geq\lambda^{2}-{(6t)}/{n}>n^{2}/4-54k^{2}$. Then it follows from Theorem 4.5 that $G$ is not $108k^{2}$-far from being bipartite. Thus, we can remove less than $108k^{2}$ edges from $G$ to obtain a bipartite subgraph. Equivalently, there exists a vertex partition $V(G)=S\cup T$ such that $e(S)+e(T)<108k^{2}$. Therefore, we get $e(S,T)=e(G)-108k^{2}>n^{2}/4-162k^{2}$, which implies $n/2-13k<|S|,|T|<n/2+13k$. Furthermore, we have $\delta(G)>n/2-56k^{2}$. Otherwise, if $d(v)\leq n/2-56k^{2}$ for some $v\in V(G)$, then $e(G\setminus\{v\})\geq n^{2}/4-54k^{2}-(n/2-56k^{2})>(n-1)^{2}/4+k^{2}$, which leads to a copy of $F_{k}$ in $G\setminus\{v\}$, a contradiction. Since $\delta(G)>n/2-56k^{2}$, we can show that $G[S]$ and $G[T]$ are $K_{1,k}$-free. Then $\Delta(G)<(n/2+13k)+k\leq n/2+14k$. ∎
Our key innovation in the above is to exploit the supersaturation-stability. The bipartition obtained from Lemma 4.7 could be served as [8, Lemma 15]. Consequently, we provide a new method to simplify so many lemmas as stated in [8] that we can get rid of the use of the triangle removal lemma. Hence we can drop the condition on $n$ being sufficiently large. In particular, Li, Lu and Peng [31] revisited the bowtie $F_{2}$ and obtained a tight bound $n\geq 7$ on the order of the extremal graph. Additionally, Lemma 4.7 can be also applied to the proof of a recent result due to Lin, Zhai and Zhao [36, Theorem 7]. Incidentally, we point out that Theorem 4.5 may be used to get better bounds for treating the extremal problems on $C_{2k+1}$-free graphs and $kC_{3}$-free graphs. Indeed, for example, if $G$ is a $C_{2k+1}$-free graph, then $3t(G)=\sum_{v\in V}e(N_{G}(v))\leq\sum_{v\in V}\mathrm{ex}(d_{G}(v),P_{2k})<kn^{2}$ by Erdős–Gallai’s theorem [5], so the supersaturation-stability in Theorem 4.5 can be applied as well.
③ In what follows, we shall present the third application by giving an alternative new proof of Theorem 1.4. Our proof is completely different from that in the original paper [52], and it relies mainly on the supersaturation-stability, while the original proof is based on the structural analysis on the entries of the Perron vector and on counting the walks of length two starting from the largest entry of Perron vector. Furthermore, our proof allows us to characterize the extremal graphs in Theorem 1.4. In other words, we can determine all the extremal graphs $G$ satisfying $\lambda(G)>\lambda(T_{n,2})$ and $t(G)=\left\lfloor{n}/{2}\right\rfloor-1$. Next, we briefly describe the main steps as below.
New proof of Theorem 1.4.
Assume that $G$ is an $n$-vertex graph with $\lambda(G)\geq\lambda(T_{n,2})$ and $G\neq T_{n,2}$. Moreover, we assume further that $G$ has the minimum number of triangles. Then $t(G)\leq\left\lfloor\frac{n}{2}\right\rfloor-1\leq\frac{n-2}{2}$. Note that $\lambda(G)\geq\lambda(T_{n,2})>\frac{n-1}{2}$. By Lemma 3.6, we get
$e(G)\geq\lambda^{2}-\frac{3t}{\lambda}>\lambda^{2}-\frac{6t}{n-1}\geq\left\lfloor\frac{n^{2}}{4}\right\rfloor-\frac{3(n-2)}{n-1}.$
Note that $e(G)$ must be an integer. Then
$e(G)\geq\left\lfloor\frac{n^{2}}{4}\right\rfloor-2.$
If $G$ is $6$-far from being bipartite, then Theorem 4.5 implies that
$t(G)\geq\frac{n}{6}\left(e(G)+6-\frac{n^{2}}{4}\right)>\frac{n}{2},$
a contradiction. Thus $G$ is not $6$-far from being bipartite. Consequently, there is a partition of the vertex set of $G$ as $V(G)=S\cup T$ such that $e(S)+e(T)<6$. Then
$e(S,T)=e(G)-e(S)-e(T)\geq e(G)-5\geq\left\lfloor\frac{n^{2}}{4}\right\rfloor-7.$
By AM-GM inequality, we get
$\left\lfloor\frac{n}{2}\right\rfloor-2\leq|S|,|T|\leq\left\lceil\frac{n}{2}\right\rceil+2.$
We say an edge is a class-edge of $G$ if the endpoints of this edge are in either $S$ or $T$. Similarly, an edge is said to be a cross-edge if it has one endpoint in $S$ and the other in $T$. Next, we claim that there is exactly one class-edge in $G$. Namely,
$e(S)+e(T)=1.$
Otherwise, suppose that $G$ has $s$ class-edges, where $2\leq s\leq 5$. Observe that each missing cross-edges between $S$ and $T$ is contained in at most $s$ triangles. Then for $n\geq 36$, we have $t(G)\geq s(\lfloor\frac{n}{2}\rfloor-2)-7s>\lfloor\frac{n}{2}\rfloor-1$, a contradiction. Thus, we conclude that $e(S)+e(T)=1$. Using this claim, we can make a slight refinement as below:
$e(S,T)=e(G)-1\geq\left\lfloor\frac{n^{2}}{4}\right\rfloor-3$
and
$\left\lfloor\frac{n}{2}\right\rfloor-1\leq|S|,|T|\leq\left\lceil\frac{n}{2}\right\rceil+1.$
Without loss of generality, we may assume that $e(S)=1$ and $e(T)=0$. Thus, $G$ is a subgraph of $K_{s,t}^{+}$ with $s\in[\frac{n}{2}-1,\frac{n}{2}+1]$, and $G$ satisfies $\lambda(G)\geq\lambda(T_{n,2})$ and $t(G)\leq\lfloor\frac{n}{2}\rfloor-1$. Finally, using a simple argument, we can compute that
$G\in\left\{K_{\frac{n}{2}+1,\frac{n}{2}-1}^{+},K_{\frac{n}{2},\frac{n}{2}}^{+|},K_{\frac{n+1}{2},\frac{n-1}{2}}^{+|}\right\}.$
For simplicity, we omit the tedious calculation, since a similar argument can be found in the remark after the proof of Theorem 2.1 in Section 5. ∎
Remark. Recall that a theorem of Erdős and Rademacher [15, 18] states that $t(G)\geq\lfloor{n}/{2}\rfloor$ provided by $e(G)>e(T_{n,2})$. At the first glance, it seems not comparable between Theorem 1.4 and Erdős–Rademacher’s theorem. In the above proof, when we have determined the spectral extremal graphs in Theorem 1.4, we can see that Theorem 1.4 actually implies the Erdős–Rademacher theorem. Indeed, as long as $G$ is a graph satisfying $e(G)>e(T_{n,2})$, we can get $\lambda(G)>\lambda(T_{n,2})$ by invoking the fact $\lambda(G)\geq 2e(G)/n$. Consequently, it follows from Theorem 1.4 that $t(G)\geq\lfloor{n}/{2}\rfloor-1$, while the graphs attaining the equality has exactly $\lfloor{n^{2}}/{4}\rfloor$ edges. Therefore, we have $t(G)\geq\lfloor{n}/{2}\rfloor$, as needed. Thus, it seems also meaningful to characterize the equality case of Theorem 1.4 in this sense.
5 Proof of Theorem 2.1
Assume that $G$ is a graph of order $n$ with $\lambda(G)\geq\lambda(T_{n,2})$ and $G\neq T_{n,2}$, we need to prove that $G$ has at least $12\lfloor n/2\rfloor-$ 1 triangular edges. Suppose on the contrary that $G$ has less than $12\lfloor n/2\rfloor-$ 1 triangular edges (This bound can be changed to $12\lfloor n/2\rfloor+$ 1 in order to adapt the proof of Theorem 2.2). Among such counterexamples, we choose $G$ as a graph with the maximum spectral radius.
Lemma 5.1.
There exists a vertex partition $V(G)=S\cup T$ such that
$e(S)+e(T)<6\sqrt{n}$
and
$e(S,T)>\frac{n^{2}}{4}-9\sqrt{n}.$
Furthermore, we have
$\frac{n}{2}-3n^{1/4}<|S|,|T|<\frac{n}{2}+3n^{1/4}.$
Proof.
Since $G$ has less than $n$ triangular edges, we know from Kruskal–Katona’s theorem (see, e.g., [5, page 305]) that $G$ has less than $\sqrt{2}n^{3/2}/3<n^{3/2}/2$ triangles. Note that $\lambda(G)\geq\lambda(T_{n,2})=\sqrt{\lfloor n^{2}/4\rfloor}>{(n-1)}/{2}$. Then Lemma 3.6 implies
$e(G)\geq\lambda^{2}-\frac{6t}{n-1}>\frac{n^{2}}{4}-3\sqrt{n}.$
We claim that $G$ is not $6\sqrt{n}$-far from being bipartite. Suppose on the contrary that $G$ is $6\sqrt{n}$-far from being bipartite. Then Theorem 4.5 implies that $G$ has at least ${n}/{6}({n^{2}}/{4}-3\sqrt{n}+6\sqrt{n}-{n^{2}}/{4})=n^{3/2}/2$ triangles, a contradiction. Therefore, $G$ is not $6\sqrt{n}$-far from being bipartite. Namely, there exists a vertex partition of $G$ as $V(G)=S\cup T$ such that
$e(S)+e(T)<6\sqrt{n}.$
Consequently, we get
$e(S,T)>e(G)-6\sqrt{n}>\frac{n^{2}}{4}-9\sqrt{n}.$
Without loss of generality, we may assume that $1\leq|S|\leq|T|$. Suppose on the contrary that $|S|\leq{n}/{2}-3n^{1/4}$. Then by $|S|+|T|=n$, we have $|T|\geq{n}/{2}+3n^{1/4}$. It follows that $e(S,T)\leq|S||T|\leq({n}/{2}-3n^{1/4})({n}/{2}+3n^{1/4})={n^{2}}/{4}-9n^{1/2}$, a contradiction. Thus, we obtain $|S|>{n}/{2}-3n^{1/4}$ and $|T|=n-|S|<{n}/{2}+3n^{1/4}$, as required. ∎
In the sequel, we assume that $V(G)=S\cup T$ is a partition with maximum cut, i.e., the bipartite subgraph $G[S,T]$ has the maximum number of edges. Furthermore, Lemma 5.1 guarantees that there exists such a partition with $e(S,T)>n^{2}/4-9\sqrt{n}$ and $e(S)+e(T)<6\sqrt{n}$. Next, we define two sets of ‘bad’ vertices of $G$. Namely, we define
$L:=\left\{v\in V(G):d(v)\leq\left(\frac{1}{2}-\frac{1}{200}\right)n\right\}.$
For a vertex $v\in V(G)$, let $d_{S}(v)=|N(v)\cap S|$ and $d_{T}(v)=|N(v)\cap T|$. We denote
$W:=\left\{v\in S:d_{S}(v)\geq\frac{n}{140}\right\}\cup\left\{v\in T:d_{T}(v)\geq\frac{n}{140}\right\}.$
First of all, we shall show that both $W$ and $L$ are small sets.
Lemma 5.2.
We have
$|L|<10.$
Proof.
Suppose that $|L|\geq 10$. Then let $L^{\prime}\subseteq L$ with $|L^{\prime}|=10$. We consider the subgraph of $G$ by deleting all vertices of $L^{\prime}$. It follows that
$e(G\setminus L^{\prime})$$\geq$ $e(G)-\sum_{v\in L^{\prime}}d(v)$
$\geq$ $\frac{n^{2}}{4}-3\sqrt{n}-10\left(\frac{1}{2}-\frac{1}{200}\right)n$
$\geq$ $\frac{(n-10)^{2}}{4}+25,$
where the last inequality holds for $n\geq 5416$. By modifying the proof of Theorem 1.1, we can see that the subgraph $G\setminus L^{\prime}$ contains more than $n+1$ triangular edges, a contradiction (In fact, a result of Füredi and Maleki [23] can indicate more triangular edges in $G\setminus L^{\prime}$). So we have $|L|<10$. ∎
Lemma 5.3.
We have
$|W|<\frac{1680}{\sqrt{n}}.$
Proof.
We denote $W_{1}=W\cap S$ and $W_{2}=W\cap T$. Then
$2e(S)=\sum_{u\in S}d_{S}(u)\geq\sum_{u\in W_{1}}d_{S}(u)\geq\frac{n}{140}|W_{1}|$
and
$2e(T)=\sum_{u\in T}d_{T}(u)\geq\sum_{u\in W_{2}}d_{T}(u)\geq\frac{n}{140}|W_{2}|.$
So we obtain
$e(S)+e(T)\geq(|W_{1}|+|W_{2}|)\frac{n}{280}=\frac{|W|n}{280}.$
On the other hand, by Lemma 5.1, we have
$e(S)+e(T)<6\sqrt{n}.$
Then we get ${|W|n}/{280}<6\sqrt{n}$, that is, $|W|<{1680}/{\sqrt{n}}$, as needed. ∎
We will also need the following inclusion-exclusion principle.
Lemma 5.4.
Let $A_{1},A_{2},\dots,A_{k}$ be $k$ finite sets. Then
$\left|\bigcap_{i=1}^{k}A_{i}\right|\geq\sum_{i=1}^{k}|A_{i}|-(k-1)\left|\bigcup_{i=1}^{k}A_{i}\right|.$
Lemma 5.5.
We have $W\subseteq L$ and $|W|\leq|L|<10$.
Proof.
We shall prove that if $u\notin L$, then $u\notin W$. We denote $L_{1}=L\cap S$ and $L_{2}=L\cap T$. Without loss of generality, we may assume that $u\in S$ and $u\notin L_{1}.$ Since $S$ and $T$ form a maximum cut in $G$, we claim that $d_{T}(u)\geq\frac{1}{2}d(u)$. Otherwise, if $d_{T}(u)<\frac{1}{2}d(u)$, then by $d(u)=d_{S}(u)+d_{T}(u)$, we have $d_{S}(u)>d_{T}(u)$. Removing the vertex $u$ from $S$ to $T$ yields a new vertex bipartition with more edges, which contradicts with the maximality of $G[S,T]$. So we must have $d_{T}(u)\geq\frac{1}{2}d(u)$. On the other hand, we have $d(u)>\left(\frac{1}{2}-\frac{1}{200}\right)n$ since $u\not\in L$. Then
$d_{T}(u)\geq\frac{1}{2}d(u)>\left(\frac{1}{4}-\frac{1}{400}\right)n.$
Recall that $|L|<10$ and $|W|<1680/\sqrt{n}$, we have $|S\setminus(W\cup L)|\approx\frac{n}{2}$. We claim that $u$ has at most $7$ neighbors in $S\setminus(W\cup L)$. Indeed, suppose on the contrary that $u$ is adjacent to $8$ vertices $u_{1},u_{2},\ldots,u_{8}$ in $S\setminus(W\cup L)$. Since $u_{i}\not\in L$, we have $d(u_{i})>\left(\frac{1}{2}-\frac{1}{200}\right)n$. Similarly, we have $d_{S}(u_{i})<\frac{n}{140}$ as $u_{i}\notin W$. So
$d_{T}(u_{i})=d(u_{i})-d_{S}(u_{i})>\left(\frac{1}{2}-\frac{1}{200}-\frac{1}{140}\right)n.$
By Lemma 5.4, we have
$\left|N_{T}(u)\cap N_{T}(u_{1})\cap\cdots\cap N_{T}(u_{8})\right|$
$\geq$ $|N_{T}(u)|+|N_{T}(u_{1})|+\cdots+|N_{T}(u_{8})|-8|T|$
$>$ $\left(\frac{1}{4}-\frac{1}{400}\right)n+\left(\frac{1}{2}-\frac{1}{200}-\frac{1}{140}\right)n\cdot 8-8\left(\frac{n}{2}+3n^{1/4}\right)$
$>$ $\frac{n}{9},$
where the last inequality holds for $n\geq 5191$. Let $B$ be the common neighbors of $u,u_{1},\ldots,u_{8}$ in $T$. Then $|B|>n/9$. Observe that for each vertex $v\in B$, the $vuu_{i}$ forms a triangle for each $1\leq i\leq 8$, so $vu,vu_{i}(1\leq i\leq 8)$ are triangular edges. That is to say, each vertex of $B$ is incident to at least $9$ triangular edges. This leads to more than $89|B|+8>n+$ 8 triangular edges, a contradiction. Therefore $u$ is adjacent to at most $7$ vertices in $S\setminus(W\cup L)$. Recall that $|L|\leq 9$ by Lemma 5.2. Hence, for $n\geq 5432$, we have
$d_{S}(u)\leq|W|+|L|+7<\frac{1680}{\sqrt{n}}+16<\frac{n}{140}.$
By definition, we get $u\notin W$. This completes the proof. ∎
Lemma 5.6.
We have $e(S\setminus L)\leq 1$ and $e(T\setminus L)\leq 1$. Consequently, there exist independent sets $I_{S}\subseteq S\setminus L$ and $I_{T}\subseteq T\setminus L$ such that
$|I_{S}|\geq|S|-10$
and
$|I_{T}|\geq|T|-10.$
Proof.
First of all, we show that $e(S\setminus L)\leq 1$ and $e(T\setminus L)\leq 1$. Suppose on the contrary that $G[S\setminus L]$ contains two edges, say $e_{1},e_{2}$. If $e_{1}$ and $e_{2}$ are intersecting, then we assume that $e_{1}=\{u_{1},u_{2}\}$ and $e_{2}=\{u_{1},u_{3}\}$. Since $u_{1},u_{2},u_{3}\notin L$, we get $d(u_{i})>\left(\frac{1}{2}-\frac{1}{200}\right)n$. By Lemma 5.5, we have $u_{i}\notin W$ and $d_{S}(u_{i})<\frac{n}{140}$. Hence $d_{T}(u_{i})=d(u_{i})-d_{S}(u_{i})>\left(\frac{1}{2}-\frac{1}{200}-\frac{1}{140}\right)n.$ By Lemma 5.4, we get
$\left|\bigcap_{i=1}^{3}N_{T}(u_{i})\right|$$\geq$ $\sum_{i=1}^{3}|N_{T}(u_{i})|-2\left|\bigcup_{i=1}^{3}N_{T}(u_{i})\right|$
$>$ $\left(\frac{1}{2}-\frac{1}{200}-\frac{1}{140}\right)n\cdot 3-2\left(\frac{n}{2}+3n^{1/4}\right)$
$>$ $\frac{n}{3},$
where the last inequality follows by $n\geq 166$. Consequently, each vertex of the common neighbors of $\{u_{1},u_{2},u_{3}\}$ leads to at least $3$ new triangular edges, so $G$ has more than $n$ triangular edges, which is a contradiction. If $e_{1}$ and $e_{2}$ are disjoint, then we denote $e_{1}=\{u_{1},u_{2}\}$ and $e_{2}=\{u_{3},u_{4}\}$. Similarly, we can see that
$\left|\bigcap_{i=1}^{4}N_{T}(u_{i})\right|$$\geq$ $\sum_{i=1}^{4}|N_{T}(u_{i})|-3\left|\bigcup_{i=1}^{4}N_{T}(u_{i})\right|$
$>$ $\left(\frac{1}{2}-\frac{1}{200}-\frac{1}{140}\right)n\cdot 4-3\left(\frac{n}{2}+3n^{1/4}\right)$
$>$ $\frac{n}{4},$
where the last inequality holds for $n\geq 159$. In this case, we can also find more than $n$ triangular edges in $G$, a contradiction. Therefore, we conclude that $e(S\setminus L)\leq 1$. Now, by deleting at most one vertex from an edge in $G[S\setminus L]$, we can obtain a large independent set. Since $|L|\leq 9$ as Lemma 5.2, there exists an independent set $I_{S}\subseteq S\setminus L$ such that $|I_{S}|\geq|S\setminus L|-1\geq|S|-10$ by Lemma 5.5. The same argument gives that there is an independent set $I_{T}\subseteq T\setminus L$ with $|I_{T}|\geq|T|-10$. ∎
Let $\mathbf{x}\in\mathbb{R}^{n}$ be an eigenvector vector corresponding to $\lambda(G)$. By the Perron–Frobenius theorem, we know that $\mathbf{x}$ has all non-negative entries. For a vertex $v\in V(G)$, we will write ${x}_{v}$ for the eigenvector entry of $\mathbf{x}$ corresponding to $v$. Let $z$ be a vertex with the maximum eigenvector entry. We may assume by scaling that ${x}_{z}=1$.
Lemma 5.7.
We have $\sum\limits_{v\in I_{T}}{x}_{v}>\frac{n}{2}-21$.
Proof.
Considering $z$-th entry of the eigenvector equation $A(G)\mathbf{x}=\lambda\mathbf{x}$, we have
$\frac{n-1}{2}<\lambda(G)=\lambda(G){x}_{z}=\sum_{v\in N(z)}{x}_{v}\leq d(z).$
Hence $z\notin L$. Without loss of generality, we may assume that $z\in S$. By Lemma 5.5, we know that $W\subseteq L$ and $|L|\leq 9$. From Lemma 5.6, we have $d_{S\setminus L}(z)\leq 1$ and
$d_{S}(z)\leq d_{S\setminus L}(z)+|L|\leq 10.$
Therefore, we get
$\lambda(G)$ $=$ $\lambda(G){x}_{z}=\sum_{v\in N_{S}(z)}{x}_{v}+\sum_{v\in N_{T}(z)}{x}_{v}$
$=$ $\sum_{v\in N_{S}(z)}{x}_{v}+\sum_{v\sim z,v\in I_{T}}{x}_{v}+\sum_{v\sim z,v\in T\setminus I_{T}}{x}_{v}$
$\leq$ $10+\sum_{v\in I_{T}}{x}_{v}+|T\setminus I_{T}|$
$\leq$ $\sum_{v\in I_{T}}{x}_{v}+20.$
Recall that $\lambda(G)\geq\lambda(T_{n,2})>\frac{n-1}{2}$. So $\sum\limits_{v\in I_{T}}{x}_{v}>\frac{n}{2}-21$, as desired. ∎
Lemma 5.8.
We have $L=\varnothing$ and $e(S)+e(T)\leq 1$.
Proof.
By way of contradiction, assume that there is a vertex $v\in L$, then $d(v)\leq(\frac{1}{2}-\frac{1}{200})n$. We define a graph $G^{+}$ with the vertex set $V(G)$ and the edge set
$E(G^{+})=E(G\setminus\{v\})\cup\{vw:w\in I_{T}\}.$
Note that adding a vertex incident with vertices in $I_{T}$ does not create any triangular edges since $I_{T}$ is an independent set. By Lemma 5.7, we have
$\lambda(G^{+})-\lambda(G)$$\geq$ $\frac{\mathbf{x}^{T}\left(A(G^{+})-A(G)\right)\mathbf{x}}{\mathbf{x}^{T}\mathbf{x}}$
$=$ $\frac{2{x}_{v}}{\mathbf{x}^{T}\mathbf{x}}\biggl{(}\sum_{w\in I_{T}}{x}_{w}-\sum_{u\in N_{G}(v)}{x}_{u}\biggr{)}$
$>$ $\frac{2{x}_{v}}{\mathbf{x}^{T}\mathbf{x}}\left(\frac{n}{2}-21-\left(\frac{1}{2}-\frac{1}{200}\right)n\right)$
$=$ $\frac{2{x}_{v}}{\mathbf{x}^{T}\mathbf{x}}\left(\frac{n}{200}-21\right)>0,$
where the last inequality holds for $n>4200$. This contradicts with the maximality of the spectral radius of $G$, so $L$ must be empty.
By Lemma 5.6, we get $e(S)\leq 1$ and $e(T)\leq 1$. Since $L=\varnothing$, then for every vertex $v\in S$, we have $d(v)>(\frac{1}{2}-\frac{1}{200})n$ and $d_{S}(v)\leq 1$. So $d_{T}(v)\geq(\frac{1}{2}-\frac{1}{200})n$. The corresponding degree condition also holds for each vertex of $T$. Suppose on the contrary that $e(S)=1$ and $e(T)=1$. Then we denote $e_{1}=\{v_{1},v_{2}\}\in E(G[S])$. Observe that for $n\geq 108$, we have
$|N_{T}(v_{1})\cap N_{T}(v_{2})|>2\left(\frac{1}{2}-\frac{1}{200}\right)n-\left(\frac{n}{2}+3n^{1/4}\right)>\frac{2n}{5}.$
Each vertex of the common neighbors of $v_{1},v_{2}$ in $T$ can yield two triangular edges. There are more than $\frac{4}{5}n$ triangular edges between $\{v_{1},v_{2}\}$ and $N_{T}(v_{1})\cap N_{T}(v_{2})$. Similarly, the edge in $G[T]$ can lead to at least $\frac{4n}{5}-4$ new triangular edges, so $G$ has more than $\frac{7}{5}n$ triangular edges. This is a contradiction. Therefore, we have $e(S)+e(T)\leq 1$, as required. ∎
The most general result is a structure theorem which asserts that any graph with larger spectral radius than $T_{n,2}$ and few triangular edges can be approximated by an almost-balanced complete bipartite graph. Just like the classical stability method, once we have proved that the extremal graph is quite close to the conjectured graph, we can show further that it must be exactly the conjectured graph.
Theorem 5.9.
If $G$ is a spectral extremal graph of order $n$ with at most $n+1$ triangular edges, then $e(G)\geq\lfloor{n^{2}}/{4}\rfloor-3$. Moreover, there exists a vertex partition $V(G)=S\cup T$ such that $e(S,T)\geq\lfloor n^{2}/4\rfloor-4$ and $\lceil{n}/{2}\rceil-2\leq|S|,|T|\leq\lfloor{n}/{2}\rfloor+2$.
Proof.
From Lemma 5.8, we have $e(S)+e(T)\leq 1$. Since any triangle contains an edge of $E(S)\cup E(T)$, the number of triangles in $G$ is bounded above by $\frac{n}{2}+3n^{1/4}$. By Lemma 3.6, we have
$e(G)\geq\lambda^{2}-\frac{6t}{n-1}>\left\lfloor\frac{n^{2}}{4}\right\rfloor-4.$
Then
$e(S,T)=e(G)-e(S)-e(T)>\frac{n^{2}}{4}-5.$
By symmetry, we may assume that $|S|\leq|T|$. Suppose on the contrary that $|S|\leq\lceil\frac{n}{2}\rceil-3$. Then $|T|=n-|S|\geq\lfloor\frac{n}{2}\rfloor+3$. If $n$ is even, then it follows that $e(S,T)\leq|S||T|\leq\left(\frac{n}{2}-3\right)\left(\frac{n}{2}+3\right)=\frac{n^{2}}{4}-9$, which contradicts with $e(S,T)\geq{n^{2}}/{4}-4$. If $n$ is odd, then $e(S,T)\leq(\frac{n+1}{2}-3)(\frac{n-1}{2}+3)=\frac{n^{2}-1}{4}-6$, a contradiction. Thus, we have
$\left\lceil\frac{n}{2}\right\rceil-2\leq|S|,|T|\leq\left\lfloor\frac{n}{2}\right\rfloor+2.$
This completes the proof. ∎
Now, we are ready to present the proof of Theorem 2.1.
Proof of Theorem 2.1.
Assume that $G$ has at most $22\lfloor n/2\rfloor-$ 2 triangular edges and $G\neq T_{n,2}$. We need to prove that $\lambda(G)<\lambda(T_{n,2})$. Furthermore, we can choose $G$ as a graph with the maximum spectral radius. Suppose on the contrary that $\lambda(G)\geq\lambda(T_{n,2})$. In what follows, we will deduce a contradiction.
First of all, we know from Theorem 1.4 that $G$ contains at least $\lfloor n/2\rfloor-1$ triangles We use Theorem 1.4 in order to avoid the complicated computations.. By Theorem 5.9, $G$ is almost complete bipartite, and we have $n/2-2\leq|S|,|T|\leq n/2+2$. If $e(S)+e(T)=0$, then $G$ is a bipartite graph with color classes $S$ and $T$. So we have $\lambda(G)\leq\sqrt{|S||T|}\leq\sqrt{\lfloor n^{2}/4\rfloor}$ since $|S|+|T|=n$. On the other hand, our assumption gives $\lambda(G)\geq\lambda(T_{n,2})=\sqrt{\lfloor n^{2}/4\rfloor}$. Therefore, it follows that $G=T_{n,2}$, a contradiction. By Lemma 5.8, we now assume that $e(S)+e(T)=1$. Next, we proceed the proof in two cases.
Case 1. Assume that $n$ is even.
Subcase 1.1. $|S|=\frac{n}{2}-2$ and $|T|=\frac{n}{2}+2$. If $e(S)=1$, then $G$ is a subgraph of $K_{\frac{n}{2}-2,\frac{n}{2}+2}^{+}$. Similarly, we get that $\lambda(K_{\frac{n}{2}-2,\frac{n}{2}+2}^{+})$ is the largest root of
$g_{1}(x)=x^{3}-x^{2}+4x-(n^{2}x)/4+n^{2}/4-n-8.$
We can check that $g_{1}(\frac{n}{2})=n-8>0$ and $g_{1}^{\prime}(x)\geq 0$ for every $x\geq\frac{n}{2}$. It follows that $\lambda(K_{\frac{n}{2}-2,\frac{n}{2}+2}^{+})<\frac{n}{2}$. If $e(T)=1$, then $G$ is a subgraph of $K_{\frac{n}{2}+2,\frac{n}{2}-2}^{+}$. By computation, we obtain that $\lambda(K_{\frac{n}{2}+2,\frac{n}{2}-2}^{+})$ is the largest root of
$g_{2}(x)=x^{3}-x^{2}+4x-(n^{2}x)/4+n^{2}/4-n.$
It is easy to verify that $g_{2}(\frac{n}{2})=n>0$ and $g_{2}^{\prime}(x)\geq 0$ for $x\geq\frac{n}{2}$. Thus, we have $\lambda(K_{\frac{n}{2}+2,\frac{n}{2}-2}^{+})<\frac{n}{2}=\lambda(T_{n,2})$, a contradiction. Apart from the direct computation, there is another way to see that $\lambda(K_{\frac{n}{2}+2,\frac{n}{2}-2}^{+})<\lambda(T_{n,2})$. Suppose on the contrary that $\lambda(K_{\frac{n}{2}+2,\frac{n}{2}-2}^{+})\geq\lambda(T_{n,2})$. Then Theorem 1.4 implies that $K_{\frac{n}{2}+2,\frac{n}{2}-2}^{+}$ contains at least $\frac{n}{2}-1$ triangles, which is a contradiction immediately.
In this subcase, we conclude that either $\lambda(G)\leq\lambda(K_{\frac{n}{2}+2,\frac{n}{2}-2}^{+})<\lambda(T_{n,2})$ or $\lambda(G)\leq\lambda(K_{\frac{n}{2}-2,\frac{n}{2}+2}^{+})<\lambda(T_{n,2})$, which contradicts with the assumption.
Subcase 1.2. $|S|=\frac{n}{2}-1$ and $|T|=\frac{n}{2}+1$. If $e(S)=1$, then $G$ is a subgraph of $K_{\frac{n}{2}-1,\frac{n}{2}+1}^{+}$. Since $G$ has at most $n-2$ triangular edges, and $K_{\frac{n}{2}-1,\frac{n}{2}+1}^{+}$ has $32|T|+1=n+$ 3 triangular edges. Therefore, we must destroy at least $5$ triangular edges from $K_{\frac{n}{2}-1,\frac{n}{2}+1}^{+}$ to obtain the graph $G$. Consequently, the deleted triangular edges are incident to at least $3$ vertices of $T$. Then $G$ has at most $|T|-3=\frac{n}{2}-2$ triangles, a contradiction. If $e(T)=1$, then $G$ is a subgraph of $K_{\frac{n}{2}+1,\frac{n}{2}-1}^{+}$. Observe that $K_{\frac{n}{2}+1,\frac{n}{2}-1}^{+}$ has $n-1$ triangular edges. We must delete at least one triangular edge of $K_{\frac{n}{2}+1,\frac{n}{2}-1}^{+}$ to obtain $G$. It follows that $G$ has at most $\frac{n}{2}-2$ triangles, a contradiction.
Subcase 1.3. $|S|=\frac{n}{2}$ and $|T|=\frac{n}{2}$. In this situation, we may assume by the symmetry that $e(S)=1$. Then $G$ is a subgraph of $K_{\frac{n}{2},\frac{n}{2}}^{+}$. Recall that $G$ has at most $n-2$ triangular edges, and $K_{\frac{n}{2},\frac{n}{2}}^{+}$ has exactly $n+1$ triangular edges. To obtain the graph $G$, we need to destroy at least $3$ triangular edges from $K_{\frac{n}{2},\frac{n}{2}}^{+}$. Consequently, $G$ has at most $\frac{n}{2}-2$ triangles, which is a contradiction.
Case 2. Suppose that $n$ is odd. In this case, by assumption, we know that $G$ contains at least $\frac{n-3}{2}$ triangles and $G$ has at most $n-3$ triangular edges.
Subcase 2.1. $|S|=\frac{n-3}{2}$ and $|T|=\frac{n+3}{2}$. If $e(S)=1$, then $G$ is a subgraph of $K_{\frac{n-3}{2},\frac{n+3}{2}}^{+}$. Notice that $K_{\frac{n-3}{2},\frac{n+3}{2}}^{+}$ has exactly $n+4$ triangular edges. To obtain the graph $G$, we need to destroy at least $7$ triangular edges. Then we need to delete some triangular edges that are incident to at least $4$ vertices of $T$, so $G$ has at most $|T|-4=\frac{n-5}{2}$ triangles, a contradiction. If $e(T)=1$, then $G$ is a subgraph of $K_{\frac{n+3}{2},\frac{n-3}{2}}^{+}$. By computation, we obtain that $\lambda(K_{\frac{n+3}{2},\frac{n-3}{2}}^{+})$ is the largest root of
$g_{3}(x)=x^{3}-x^{2}+(9x)/4-(xn^{2})/4+n^{2}/4-n+3/4.$
It is easy to check that $g_{3}(\frac{1}{2}\sqrt{n^{2}-1})=1-n+\sqrt{n^{2}-1}>0$. Moreover, we have $g_{3}^{\prime}(x)=3x^{2}-2x+9/4-n^{2}/4$. We can verify that $g_{3}^{\prime}(x)>0$ for any $x>\frac{1}{2}\sqrt{n^{2}-1}$, which yields $\lambda(G)\leq\lambda(K_{\frac{n+3}{2},\frac{n-3}{2}}^{+})<\frac{1}{2}\sqrt{n^{2}-1}=\lambda(T_{n,2})$, a contradiction.
Subcase 2.2. $|S|=\frac{n-1}{2}$ and $|T|=\frac{n+1}{2}$. If $e(S)=1$, then $G$ is a subgraph of $K_{\frac{n-1}{2},\frac{n+1}{2}}^{+}$. Since $K_{\frac{n-1}{2},\frac{n+1}{2}}^{+}$ has exactly $n+2$ triangular edges, we must destroy at least $5$ triangular edges to obtain the graph $G$. So the deleted triangular edges are incident to at least $3$ vertices of $T$, and $G$ contains at most $|T|-3=\frac{n-5}{2}$ triangles, a contradiction. If $e(T)=1$, then $G$ is a subgraph of $K_{\frac{n+1}{2},\frac{n-1}{2}}^{+}$. As $K_{\frac{n+1}{2},\frac{n-1}{2}}^{+}$ has $n$ triangular edges, we need to destroy at least $3$ triangular edges to produce $G$. In this process, at least two triangles of $K_{\frac{n+1}{2},\frac{n-1}{2}}^{+}$ are removed, so $G$ has at most $|S|-2=\frac{n-5}{2}$ triangles, which is a contradiction. ∎
Remark. In the above proof, we can determine the extremal graphs $G$ in the sense that $\lambda(G)\geq\lambda(T_{n,2})$, $G\neq T_{n,2}$ and $G$ has exactly $12\lfloor\frac{n}{2}\rfloor-$ 1 triangular edges. Indeed, we next give the sketch without the details.
In Subcase 1.1, it was proved that $\lambda(G)<\lambda(T_{n,2})$, a contradiction.
In Subcase 1.2, as we know, $G$ has exactly $n-1$ triangular edges. If $e(S)=1$, then $G$ is obtained from $K_{\frac{n}{2}-1,\frac{n}{2}+1}^{+}$ by deleting at least two triangular edges that incident to two vertices of $T$. In this deletion, we destroy four triangular edges of $K_{\frac{n}{2}-1,\frac{n}{2}+1}^{+}$. More precisely, let $\{u,v\}$ be the unique edge of $G[S]$. Then we can delete two triangular edges intersecting in $u$, or delete two disjoint triangular edges incident to $u$ and $v$, respectively. In each case, we can compute that the resulting graphs have spectral radius less than $\lambda(T_{n,2})$. If $e(T)=1$, then $G$ is a subgraph of $K_{\frac{n}{2}+1,\frac{n}{2}-1}^{+}$. Note that we can not delete any triangular edges to obtain $G$. Moreover, we can verify that the deletion of a non-triangular edge leads to a graph with spectral radius less than $\lambda(T_{n,2})$. So we have $G=K_{\frac{n}{2}+1,\frac{n}{2}-1}^{+}$. In addition, Lemma 3.1 gives $\lambda(K_{\frac{n}{2}+1,\frac{n}{2}-1}^{+})>\lambda(T_{n,2})$. Thus $K_{\frac{n}{2}+1,\frac{n}{2}-1}^{+}$ is one of the extremal graphs.
In Subcase 1.3, $G$ is obtained from $K_{\frac{n}{2},\frac{n}{2}}^{+}$ by deleting at least one triangular edge. So $G$ is a subgraph of $K_{\frac{n}{2},\frac{n}{2}}^{+|}$. By calculation, deleting any edge from $K_{\frac{n}{2},\frac{n}{2}}^{+|}$ yields a graph with spectral radius less than $\lambda(T_{n,2})$. Then we must have $G=K_{\frac{n}{2},\frac{n}{2}}^{+|}$. From Lemma 3.2, we get $\lambda(K_{\frac{n}{2},\frac{n}{2}}^{+|})>\lambda(T_{n,2})$. So $K_{\frac{n}{2},\frac{n}{2}}^{+\,|}$ is the second extremal graph.
In Subcase 2.1, $G$ has exactly $n-2$ triangular edges. If $e(S)=1$, then $G$ is a subgraph of $K_{\frac{n-3}{2},\frac{n+3}{2}}^{+}$ by deleting at least three triangular edges that incident to three vertices of $T$. For example, let $\{u,v\}$ be the unique edge of $G[S]$. We can delete three triangular edges intersecting in $u$, or we delete two triangular edges incident to $u$, and one triangular edge incident to $v$. In the two cases, the resulting subgraphs have spectral radius less than $\lambda(T_{n,2})$. If $e(T)=1$, then $G$ is a subgraph of $K_{\frac{n+3}{2},\frac{n-3}{2}}^{+}$. In the previous proof, we have shown that $\lambda(G)<\lambda(T_{n,2})$, a contradiction.
In Subcase 2.2, if $e(S)=1$, then $G$ is a subgraph of $K_{\frac{n-1}{2},\frac{n+1}{2}}^{+}$. Note that $K_{\frac{n-1}{2},\frac{n+1}{2}}^{+}$ contains $n+2$ triangular edges. We need to delete at least two triangular edges incident to two vertices of $T$. Let $\{u,v\}$ be the unique edge of $G[S]$. Then $G$ can be obtained by deleting two triangular edges intersecting in $u$, or deleting two disjoint triangular edges incident to $u$ and $v$, respectively. In both cases, we can check that the resulting graphs have spectral radius less than $\lambda(T_{n,2})$. If $e(T)=1$, then $G$ is a subgraph of $K_{\frac{n+1}{2},\frac{n-1}{2}}^{+\,|}$. We can calculate that any proper subgraph has spectral radius less than $\lambda(T_{n,2})$. Moreover, Lemma 3.3 tells us that $\lambda(K_{\frac{n+1}{2},\frac{n-1}{2}}^{+\,|})>\lambda(T_{n,2})$, so $K_{\frac{n+1}{2},\frac{n-1}{2}}^{+\,|}$ is the third extremal graph.
6 Proof of Theorem 2.2
Using a similar argument, we can prove Theorem 2.2.
Proof of Theorem 2.2.
Let $G$ be an $n$-vertex graph with $\lambda(G)\geq\lambda(K_{\lceil\frac{n}{2}\rceil,\lfloor\frac{n}{2}\rfloor}^{+})$ and $G$ has at most $12\lfloor n/2\rfloor+$ 1 triangular edges. We shall show that $G=K_{\lceil\frac{n}{2}\rceil,\lfloor\frac{n}{2}\rfloor}^{+}$. First of all, we know from Theorem 5.9 that $G$ is an almost balanced complete bipartite graph. More precisely, we have $e(G)\geq\lfloor{n^{2}}/{4}\rfloor-3$, and $G$ admits a partition $V(G)=S\cup T$ such that $e(S,T)\geq\lfloor n^{2}/4\rfloor-4$ and $n/2-2\leq|S|,|T|\leq n/2+2$. If $e(S)+e(T)=0$, then $G$ is a bipartite graph with color classes $S$ and $T$. Consequently, we get $\lambda(G)\leq\lambda(T_{n,2})<\lambda(K_{\lceil\frac{n}{2}\rceil,\lfloor\frac{n}{2}\rfloor}^{+})$, which contradicts with the assumption. By Lemma 5.8, we have $e(S)+e(T)=1$. In what follows, we proceed the proof in two cases.
Case 1. Assume that $n$ is even.
Subcase 1.1. $|S|=\frac{n}{2}-2$ and $|T|=\frac{n}{2}+2$. If $e(T)=1$, then $G$ is a subgraph of $K_{\frac{n}{2}+2,\frac{n}{2}-2}^{+}$. In the proof of Theorem 2.1 for Subcase 1.1, we have shown that $\lambda(K_{\frac{n}{2}+2,\frac{n}{2}-2}^{+})<\frac{n}{2}$, a contradiction. If $e(S)=1$, then $G$ is a subgraph of $K_{\frac{n}{2}-2,\frac{n}{2}+2}^{+}$. We also showed that $\lambda(G)\leq\lambda(K_{\frac{n}{2}-2,\frac{n}{2}+2}^{+})<\frac{n}{2}$, which contradicts with the assumption.
Subcase 1.2. $|S|=\frac{n}{2}-1$ and $|T|=\frac{n}{2}+1$. If $e(S)=1$, then $G$ is a subgraph of $K_{\frac{n}{2}-1,\frac{n}{2}+1}^{+}$. Similarly, we can show by contradiction that
$\lambda(K_{\frac{n}{2}-1,\frac{n}{2}+1}^{+})<\lambda(K_{\frac{n}{2},\frac{n}{2}}^{+}).$
Indeed, since $\lambda(K_{\frac{n}{2}-1,\frac{n}{2}+1}^{+})$ is the largest root of
$h_{1}(x)=-3-n+n^{2}/4+x-(n^{2}x)/4-x^{2}+x^{3}.$
Recall in Lemma 3.4 that $\lambda(K_{\frac{n}{2},\frac{n}{2}}^{+})$ is the largest root of
$f(x)=-n+n^{2}/4-(n^{2}x)/4-x^{2}+x^{3}.$
Observe that $h_{1}(x)-f(x)=x-3>0$ for every $x>3$. Then we have $h_{1}(x)>f(x)\geq 0$ for any $x\geq\lambda(K_{\frac{n}{2},\frac{n}{2}}^{+})$, which implies $\lambda(K_{\frac{n}{2}-1,\frac{n}{2}+1}^{+})<\lambda(K_{\frac{n}{2},\frac{n}{2}}^{+})$, as needed.
If $e(T)=1$, then $G$ is a subgraph of $K_{\frac{n}{2}+1,\frac{n}{2}-1}^{+}$. We can prove that
$\lambda(K_{\frac{n}{2}+1,\frac{n}{2}-1}^{+})<\lambda(K_{\frac{n}{2},\frac{n}{2}}^{+}).$
Indeed, since $\lambda(K_{\frac{n}{2}+1,\frac{n}{2}-1}^{+})$ is the largest root of
$h_{2}(x)=1-n+n^{2}/4+x-(n^{2}x)/4-x^{2}+x^{3},$
and $h_{2}(x)>f(x)$ for any $x>0$, which yields $\lambda(K_{\frac{n}{2}+1,\frac{n}{2}-1}^{+})<\lambda(K_{\frac{n}{2},\frac{n}{2}}^{+})$.
Subcase 1.3. $|S|=\frac{n}{2}$ and $|T|=\frac{n}{2}$. By the symmetry, we may assume that $e(S)=1$. Then $G$ is a subgraph of $K_{\frac{n}{2},\frac{n}{2}}^{+}$. Since $\lambda(G)\geq\lambda(K_{\frac{n}{2},\frac{n}{2}}^{+})$, we get $G=K_{\frac{n}{2},\frac{n}{2}}^{+}$, which is the desired extremal graph.
Case 2. Suppose that $n$ is odd.
Subcase 2.1. $|S|=\frac{n-3}{2}$ and $|T|=\frac{n+3}{2}$. If $e(S)=1$, then $G$ is a subgraph of $K_{\frac{n-3}{2},\frac{n+3}{2}}^{+}$. By calculation, we obtain that $\lambda(K_{\frac{n-3}{2},\frac{n+3}{2}}^{+})$ is the largest root of
$h_{3}(x)=-(21/4)-n+n^{2}/4+(9x)/4-(n^{2}x)/4-x^{2}+x^{3}.$
By Lemma 3.4, we know that $\lambda(K_{\frac{n+1}{2},\frac{n-1}{2}}^{+})$ is the largest root of
$g(x)=3/4-n+n^{2}/4+x/4-(n^{2}x)/4-x^{2}+x^{3}.$
Since $h_{3}(x)-g(x)=2x-6$, we get $h_{3}(x)>g(x)$ for any $x>3$, so it follows that $\lambda(K_{\frac{n-3}{2},\frac{n+3}{2}}^{+})<\lambda(K_{\frac{n+1}{2},\frac{n-1}{2}}^{+})$, a contradiction.
If $e(T)=1$, then $G$ is a subgraph of $K_{\frac{n+3}{2},\frac{n-3}{2}}^{+}$. By computation, we obtain that $\lambda(K_{\frac{n+3}{2},\frac{n-3}{2}}^{+})$ is the largest root of
$h_{4}(x)=3/4-n+n^{2}/4+(9x)/4-(n^{2}x)/4-x^{2}+x^{3}.$
It is easy to check that $h_{4}(x)>g(x)$ for any $x>0$. So we have $\lambda(K_{\frac{n+3}{2},\frac{n-3}{2}}^{+})<\lambda(K_{\frac{n+1}{2},\frac{n-1}{2}}^{+})$, which contradicts with the assumption on $G$.
Subcase 2.2. $|S|=\frac{n-1}{2}$ and $|T|=\frac{n+1}{2}$. If $e(S)=1$, then $G$ is a subgraph of $K_{\frac{n-1}{2},\frac{n+1}{2}}^{+}$. Since $G$ has at most $n$ triangular edges and $K_{\frac{n-1}{2},\frac{n+1}{2}}^{+}$ has exactly $n+2$ triangular edges, we must destroy at least two triangular edges to obtain the subgraph $G$. Let $K_{\frac{n-1}{2},\frac{n+1}{2}}^{+\,|}$ be the graph obtained from $K_{\frac{n-1}{2},\frac{n+1}{2}}^{+}$ by deleting an edge between $S$ and $T$ such that this edge is incident to the unique edge of $G[S]$. Furthermore, it follows that $G$ is a subgraph of $K_{\frac{n-1}{2},\frac{n+1}{2}}^{+\,|}$. In this case, we can show that
$\lambda(K_{\frac{n-1}{2},\frac{n+1}{2}}^{+\,|})<\frac{n}{2}<\lambda(K_{\frac{n+1}{2},\frac{n-1}{2}}^{+}).$
Indeed, since the spectral radius of $K_{\frac{n-1}{2},\frac{n+1}{2}}^{+\,|}$ is the largest root of
$h_{5}(x)=-(x/2)-2nx+(n^{2}x)/2+x^{2}-nx^{2}+x^{3}/4-(n^{2}x^{3})/4+x^{5},$
and $h_{5}(\frac{n}{2})=(-8n-24n^{2}+n^{3})/32>0$ for $n\geq 25$. Moreover, we can check that $h_{5}^{\prime}(x)>0$ for every $x\geq\frac{n}{2}$. So it yields $\lambda(K_{\frac{n-1}{2},\frac{n+1}{2}}^{+\,|})<\frac{n}{2}<\lambda(K_{\frac{n+1}{2},\frac{n-1}{2}}^{+})$ by Lemma 3.4. This is a contradiction.
If $e(T)=1$, then $G$ is a subgraph of $K_{\frac{n+1}{2},\frac{n-1}{2}}^{+}$. The assumption asserts that $\lambda(G)\geq\lambda(K_{\frac{n+1}{2},\frac{n-1}{2}}^{+})$, so we get $G=K_{\frac{n+1}{2},\frac{n-1}{2}}^{+}$, which is the expected extremal graph. ∎
7 Concluding remarks
In this paper, we have bounded the minimum number of triangular edges of a graph in terms of the spectral radius, and we have established a spectral Erdős–Faudree–Rousseau theorem. The main ideas in our proof attribute to the supersaturation-stability (Theorem 4.5) and some additional spectral techniques. Particularly, we stated in Subsection 4.3 that Theorem 4.5 can deduce a conjecture of Erdős involving the booksize of a graph. It is worth mentioning that an interesting spectral problem of Zhai and Lin [66, Problem 1.2] conjectured that every $n$-vertex graph $G$ with $\lambda(G)>\lambda(T_{n,2})$ has booksize greater than $n/6$ as well. To solve this problem, it is sufficient to show a spectral version of Theorem 4.5. To be formal, we propose the following conjecture.
Conjecture 7.1.
If $G$ is $t$-far from being bipartite, then
$t(G)\geq\frac{n}{6}\Bigl{(}\lambda(G)+t-\lambda(T_{n,2})\Bigr{)}.$
Remark. Conjecture 7.1 is stronger than the aforementioned problem of Zhai and Lin. Indeed, if $G$ is an $n$-vertex graph with $\lambda(G)>\lambda(T_{n,2})$, and $G$ has exactly $t$ triangles, then assuming Conjecture 7.1, we know that $G$ is not ${6t}/{n}$-far from being bipartite. So we can remove less than ${6t}/{n}$ edges from $G$ to destroy all $t$ triangles. Thus, one of these edges is contained in more than $n/6$ triangles, as expected.
A well-known result of Nosal [44] (see, e.g., [45, 51]) implies that if $G$ is a graph with $m$ edges and $\lambda(G)>\sqrt{m}$, then it contains a triangle. In 2023, Ning and Zhai [52] proved a counting result, which asserts that if $\lambda(G)\geq\sqrt{m}$, then $G$ has at least $\lfloor\frac{\sqrt{m}-1}{2}\rfloor$ triangles, unless $G$ is a complete bipartite graph (possibly with some isolated vertices). Inspired by this result and Theorem 3.7, we propose the following problem.
Conjecture 7.2.
If $G$ is a graph with $m$ edges and
$\lambda(G)\geq\sqrt{m},$
then $G$ has at least $\sqrt{m}$ triangular edges, unless $G$ is a complete bipartite graph.
As mentioned in the introduction, we shall conclude some spectral extremal problems concerning the minimum number of edges that occur in cliques or odd cycles. Motivated by the study on the minimum number of triangular edges among graphs with $n$ vertices and $m\geq n^{2}/4+q$ edges, we firstly propose the following conjecture, which provides a spectral version of Theorem 1.3.
Conjecture 7.3.
For any graph $G$ on $n$ vertices, there exists an $n$-vertex graph $H=G(a,b,c)$ for some integers $a,b,c$ such that $\lambda(H)\geq\lambda(G)$ and $|\mathbf{NT}(H)|\geq|\mathbf{NT}(G)|$.
Recall that $T_{n,r}$ is the $r$-partite Turán graph on $n$ vertices. The famous Turán theorem [5, p. 294] states that an $n$-vertex graph $G$ with $e(G)\geq e(T_{n,r})$ has a copy of $K_{r+1}$, unless $G=T_{n,r}$. Correspondingly, Nikiforov [46] showed that if $\lambda(G)\geq\lambda(T_{n,r})$, then $G$ contains a copy of $K_{r+1}$, unless $G=T_{n,r}$. So it is natural to consider the following extension by minimizing the number of edges that occur in $K_{r+1}$.
Problem 7.4.
Suppose that $r\geq 3$ and $G$ is an $n$-vertex graph with $\lambda(G)>\lambda(T_{n,r})$. What is the smallest number of edges of $G$ that are contained in $K_{r+1}$?
Remark. Inspired by Conjecture 1.2 and Theorem 1.3, we believe intuitively that the spectral extremal graphs in Problem 7.4 are possibly analogues of graphs of the form $G(a,b,c)$, i.e., they are perhaps constructed from a complete $r$-partite graph of order $n$ by adding an almost complete graph to one of the parts.
Apart from the number of triangular edges, Erdős, Faudree and Rousseau [19] also considered the analogous problems for longer odd cycles in graph of order $n$ with more than $\lfloor n^{2}/4\rfloor$ edges. They proved that for any $k\geq 2$, every graph on $n$ vertices with $\lfloor n^{2}/4\rfloor+1$ edges contains at least $\frac{11}{144}n^{2}-O(n)$ edges that are contained in an odd cycle $C_{2k+1}$. It turns out that the case $k\geq 2$ is quite different from the triangle case. Furthermore, Erdős, Faudree and Rousseau [19] made a stronger conjecture, which asserts that every such graphs contains at least $\frac{2}{9}n^{2}-O(n)$ edges that occur in $C_{2k+1}$. We remark that adding an extra edge into the complete balanced bipartite graph is not optimal. However, Füredi and Maleki [23] showed that this conjecture is false for $k=2$, and they constructed $n$-vertex graphs with $\lfloor n^{2}/4\rfloor+1$ edges and with only $\frac{2+\sqrt{2}}{16}n^{2}+O(n)\approx 0.213n^{2}$ edges in $C_{5}$. In 2019, Grzesik, Hu and Volec [25] obtained asymptotically sharp bounds for the smallest possible number of edges in $C_{2k+1}$. By using Razborov’s flag algebras method, they proved that if $G$ is an $n$-vertex graph with $\lfloor n^{2}/4\rfloor+1$ edges, then it contains at least $\frac{2+\sqrt{2}}{16}n^{2}-O(n^{15/8})$ edges that occur in $C_{5}$, and for $k\geq 3$, it contains at least $\frac{2}{9}n^{2}-O(n)$ edges in $C_{2k+1}$. Motivated by the aforementioned results, we may propose the following spectral problem.
Problem 7.5.
Let $G$ be a graph of order $n$ with $\lambda(G)>\lambda(T_{n,2})$. For each $k\geq 2$, what is the smallest number of edges of $G$ that occur in $C_{2k+1}$?
Declaration of competing interest
The authors declare that they have no conflicts of interest to this work.
Data availability
No data was used for the research described in the article.
Acknowledgements
The authors would like to thank Dr. Xiaocong He and Dr. Loujun Yu for carefully reading an early manuscript of this paper. Yongtao Li was supported by the Postdoctoral Fellowship Program of CPSF (No. GZC20233196), Lihua Feng was supported by the National Natural Science Foundation of China grant (No. 12271527) and the National Natural Science Foundation of Hunan Province (Nos. 2020JJ4675 and 2021JJ40707), and Yuejian Peng was supported by the National Natural Science Foundation of China grant (No. 11931002).